# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v1

@@ 2 Literature Review Methodology
To ensure a comprehensive review of the literature on trust, risk, and security in Agentic AI systems, we adopted a structured methodology inspired by best practices in systematic reviews  boland2017doing ; keele2007guidelines . This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
Research Objectives
The review was guided by the following questions:
  • RQ1: What are the key trust, risk, and security challenges posed by LLM-enabled Agentic AI systems?
@@
  • RQ3: How do existing approaches map onto the TRiSM pillars: explainability, model operations, security, and privacy?
@@
Data Sources and Search Strategy
We searched the following major digital libraries: IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar, covering publications from January 2020 to May 2025. The search was executed using Boolean combinations of keywords relevant to Agentic AI and TRiSM, including:
"agentic AI" OR multi-agent systems "(AMAS)" OR "multi-agent LLMs" OR "AI agents" OR "autonomous agents" OR "intelligent agents") AND ("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR "governance" OR "oversight" OR "compliance" OR "explainability" OR "interpretability" OR "transparency" OR "privacy" OR "data protection")
In total, 150 unique papers were initially retrieved.
Inclusion and Exclusion Criteria
@@
    • Studies that explicitly discuss agentic, multi-agent, or LLM-powered AI systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
    • Peer-reviewed articles, preprints, or whitepapers from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
@@
After screening titles and abstracts, we shortlisted 250 papers. A subsequent full-text review narrowed the selection to 180 primary studies that directly addressed one or more TRiSM pillars. In addition to peer-reviewed literature, we also incorporated relevant white papers, technical blogs, and practitioner reports to contextualize and map the emerging landscape of Agentic AI within the TRiSM framework.
Quality Assurance
To ensure the rigor and reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines kitchenham2004procedures ; keele2007guidelines . Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems? (2) Are the methods, experiments, or architectural designs well-documented and reproducible? (3) Does the paper substantively engage with at least one of the trust, risk, security, explainability, or privacy aspects? (4) Does the study offer empirical, theoretical, or normative contributions relevant to Agentic AI governance? Each criterion was rated as low, medium, or high. Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only. Quality ratings were assigned independently by two reviewers, with discrepancies resolved through discussion.
Figure 1: Taxonomy of Agentic AI systems presented in this review. It includes five major sections: Fundamentals of Agentic AI Systems (Agentic AI vs. Traditional AI Agents, LLM-Based Architecture), The TRISM Framework (Governance, Explainability, ModelOps, Application Security, Privacy), Threats and Risks (Unique Threat Vectors, Taxonomy of Risks, Case Studies), Trust and Explainability, and Evaluation Metrics (including Human-Centered and Composite Metrics).
@@ 3 Fundamentals of Agentic AI Systems
The distinction between traditional AI agents and Agentic AI systems represents a paradigm shift in autonomous system design. Traditional agents operate through predefined rules muscettola1998remote , heuristic workflows hannebauer1999formal ; cranshaw2017calendar , or deterministic logic govaerts2010development , excelling in narrow, bounded environments. In contrast, Agentic AI systems leverage foundation models (LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities. Agentic AI systems fundamentally redefine autonomy through three core innovations: 1) Multi-agent coordination where specialized agents (e.g., planners, verifiers) collaborate via structured protocols qian2023chatdev ; hong2023metagpt 2) Persistent context through memory architectures that maintain state across workflows borghoff2025human 3) Dynamic meta-orchestration that delegates tasks and resolves conflicts schneider2025generative We present a comparison of AI agents and Agentic AI in Table 2.
This architecture enables longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches de2025open . Systems like AutoGen wu2023autogen demonstrate how LLM-backed agents collectively decompose complex problems through multi-step reasoning qiao2024agent , marking a qualitative leap from reactive agents to proactive problem-solving systems.
@@
Autonomy Model × Reactive execution via scripts/rules Goal-driven planning & adaptation
$\bullet$ Fixed action sequences $\bullet$ Recursive self-improvement
Cognitive Foundation × Symbolic logic & FSMs Foundation models (LLMs/LIMs)
$\bullet$ Hand-coded KBs $\bullet$ Emergent reasoning
@@
Reasoning Approach × Deterministic/single-step Multi-step & contextual
$\bullet$ Rule-based inference $\bullet$ Chain-of-Thought (CoT)
Collaboration × Isolated execution Role-specialized coordination
@@
Temporal Context × Episodic & stateless Persistent memory
$\bullet$ Session-bound $\bullet$ VectorDB/LTM
Orchestration × Hard-coded pipelines Dynamic meta-agents
$\bullet$ Sequential workflows $\bullet$ Conflict resolution
Tool Utilization × Static integrations Planned API invocation
@@
Context Awareness × Bounded inputs Memory-augmented
$\bullet$ Limited context $\bullet$ RAG architectures
Scalability × Domain-specific Cross-domain transfer
$\bullet$ High maintenance $\bullet$ Modular composition
Exemplars $\triangleright$ MYCIN  $\triangleright$ ELIZA  $\triangleright$ Finite-state bots $\triangleright$ AutoGen  $\triangleright$ ChatDev  $\triangleright$ MetaGPT  $\triangleright$ AgentVerse
@@ 3.2 LLM-Based Agentic AI System Architecture
Agentic Multi-Agent System (AMAS) represent an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks. As illustrated in Figure 2, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem. At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation wang2025comprehensive . These agents access a Shared Toolchain Interface speight2025build to execute code, perform searches, or interact with domain-specific APIs.
Communication and coordination are facilitated through a Communication Middleware weyns2009agent , allowing agents to share goals, observations, or intermediate results. A Task Manager or Orchestrator langgraph2024 governs high-level planning, delegating subtasks to agents based on their roles or specializations. Agents can read from and write to a World Model or Shared Memory guo2024largelanguagemodelbased , which stores contextual knowledge, system state, or evolving task data. Human oversight is supported through a Human-in-the-Loop Interface cranshaw2017calendar , enabling users to prompt, correct, or halt agent behavior.
To ensure accountability, a Trust and Audit Module monitors agent actions, logs tool usage, and generates behavioral traces kaur2024building . However, this modular and distributed structure introduces significant TRiSM charli_trism_2024 challenges. With multiple autonomous agents accessing external resources, the Security Gateway becomes critical for enforcing access controls, authentication, and sandboxing zhong2023towards .
Likewise, a dedicated Privacy Management Layer is essential to prevent leakage of sensitive or personally identifiable information feretzakis2024privacy , especially when data traverses multiple agents or tools. Finally, an Explainability Interface must provide interpretable rationales for multi-agent decisions, supporting transparency and trust calibration singh2025explainable . Together, these architectural elements make AMAS powerful yet complex, raising unique and urgent questions about how to ensure their trustworthiness, mitigate systemic risks, and secure them against adversarial behaviors.
@@
Figure 2: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms). The modular design highlights adaptive intelligence and inter- agent synerg
Below, we discuss the architecture of AMAS.
@@
Language Model Core (Agent Brain). At the center of an Agentic AI system lies a LLM serving as the primary decision-making controller or “brain” Saravia_LLM_Agents_2024 . The core LLM is initialized with a user goal and a structured agent prompt (defining its role, capabilities, and tool access). It then generates step-by-step decisions or actions, interpreting instructions, producing reasoning traces, and selecting next steps in either natural language or structured action formats. In many agent frameworks, such as significant_gravitas_autogpt_2023 , Baby AGI nakajima_babyagi_archive_2024 and GPT Engineer osika_gpt_engineer_2023 , the LLM governs the full control loop, orchestrating the overall system behavior.
@@
Planning and Reasoning Module. To handle complex goals, an explicit planning mechanism decomposes tasks into manageable sub-goals. This can be done internally via chain-of-thought (CoT) or tree-of-thoughts prompting Saravia_Prompt_Engineering_Guide_Web_2024 , where the model performs intermediate reasoning before arriving at a final decision. Some implementations employ external planning systems by translating goals into structured planning languages and using classical planners for long-term decision-making. Planning is often interleaved with execution and feedback: the agent refines its plan based on outcomes, alternating between reasoning, acting, and integrating observations. Techniques like ReAct yao2023reactsynergizingreasoningacting exemplify this loop by guiding the agent through repeated reasoning–action–observation cycles, improving performance on complex tasks.
@@
Memory Module. Agentic AI systems integrate memory to maintain context across iterations. This includes short-term memory (recent interactions held within the prompt context) and long-term memory (accumulated knowledge or experiences) yang2024multi . Long-term memory is often implemented using vector databases, where key facts or past events are stored and retrieved by similarity search. By reintegrating past data into the LLM’s prompt, the agent can recall relevant information across sessions, avoid repetition, and support coherent long-term planning. Effective memory management enables adaptive, learning-driven behavior.
Tool-Use Interface. To extend its capabilities beyond text generation, the agent is equipped with a tool-use interface schick2023toolformerlanguagemodelsteach . This layer allows invocation of external tools such as web search, APIs, code interpreters, or databases. The available tools are specified in the agent prompt with command schemas. When the LLM determines a tool is needed, it emits a structured command, which is executed externally. The result is fed back into the LLM as a new observation. This mechanism enables the agent to access real-time information, perform computations, and interact with external systems dynamically. Tool-use frameworks like MRKL karpas2022mrklsystemsmodularneurosymbolic illustrate this design by routing queries to different expert modules (symbolic or learned) with an LLM as the router. Similarly, approaches like Toolformer schick2023toolformerlanguagemodelsteach train the model to insert API calls in its text generation.
Perception and Environment Interface. For agents interacting with dynamic environments, such as web interfaces, simulated worlds, or physical systems, an observation - action interface is essential liu2025agent . Perception modules translate raw inputs (e.g., sensor data, images, or textual states) into representations the LLM can process. Conversely, the agent’s chosen actions are executed within the environment and the resulting state changes are returned to the agent as observations. This loop supports sense-plan-act cycles that continue until the task is completed or halted. In robotic or multimodal settings, the interface may include additional sensory models (e.g., vision transformers), but the core flow remains consistent.
Integration and Autonomy. These modules together form a closed-loop architecture. The LLM plans and reasons over tasks, guided by memory and tools, and interfaces with the environment to execute decisions and observe results acharya2025agentic . Each iteration enriches the agent’s context, enabling it to self-prompt, generate subtasks, assess progress, and adapt strategies over time. This integrated design empowers agentic systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
Table  3 presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice. For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory, while Voyager couples the same core with an embodied Minecraft API that serves as both tool layer and environment interface. We comparing current Agentic AI implementations and identifying open design trade-offs in this table.
@@
Table 3: Representative LLM-Based Agentic AI Frameworks
@@
AutoGPT yang2023auto GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
BabyAGI nakajima_babyagi_archive_2024 GPT-3.5/4 Task queue In-memory Web search Minimal task generator
@@
GPT Engineer gpt-engineer GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
@@
LangGraph wang2024agent Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
@@
AutoGen wu2023autogen GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
@@
MRKL karpas2022mrklsystemsmodularneurosymbolic LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
@@
Reflexion shinn2023reflexionlanguageagentsverbal GPT-3.5/4 Retry–reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
@@
MetaGPT hong2024metagpt GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
@@
Voyager wang2023voyageropenendedembodiedagent GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
@@
WebVoyager he2024webvoyager LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
@@
HuggingGPT shen2023hugginggpt ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
@@
CAMEL li2023camel GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
@@
ChatDev qian2024chatdevcommunicativeagentssoftware GPT-4 Chat chain File repo Unix tools Simulated software development workflow
@@
CrewAI crewAI Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
@@
AgentVerse chen2023agentverse Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
@@
OpenAgents xie2023openagentsopenplatformlanguage Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
SuperAGI superagi GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use

### v2

@@ 3 Literature Review Methodology
To provide a comprehensive review of the literature on trust, risk, and security in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [41, 42]. While this is not a formal systematic review, we followed a transparent and reproducible process to ensure coverage and rigor. This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
(a) Agentic AI papers per year (2022–2025).
(b) TRiSM pillars papers by topic.
Figure 2: (a) Annual trend in Agentic AI research, showing a sharp rise post-2023. (b) Distribution of TRiSM-related publications across governance and safety topics.
3.1 Research Objectives
The review was guided by the following four specific questions:
  • RQ1: What are the key trust, risk, and security challenges posed by LLM-enabled Agentic AI systems? How do current approaches establish and measure the trustworthiness of Agentic AI outputs and decisions?
@@ 3.1 Research Objectives
  • RQ3: How do existing approaches map onto the TRiSM pillars: explainability (including trustworthiness), model operations, security, and privacy?
@@
3.2 Data Sources and Search Strategy
To conduct a thorough literature review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search encompassed publications from January 2022 to May 2025, with select seminal works prior to 2022 included to provide foundational context. We utilized Boolean search queries with an expanded set of keywords to capture the interdisciplinary intersection of LLM-based multi-agent systems and TRiSM considerations. The search terms were structured as follows:
(“ Agentic AI” OR “multi-agent systems” OR “MAS” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
AND (“trust” OR “trustworthiness” OR “risk” OR “security” OR “safety” OR “governance” OR “oversight” OR “compliance” OR “explainability” OR “interpretability” OR “transparency” OR “privacy” OR “data protection” OR “robustness” OR “accountability” OR “ethical AI” OR “fairness” OR “adversarial attacks” OR “regulatory compliance” OR “bias mitigation” OR “system reliability”)
This comprehensive search strategy ensured the identification of relevant studies addressing the integration of TRiSM principles with LLM-based multi-agent systems, particularly in high-stakes domains such as healthcare, science, and finance.
3.3 Inclusion and Exclusion Criteria
@@
    • Studies that explicitly discuss Agentic AI, Multi-Agents, or LLM-powered AI systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
    • Peer-reviewed articles, preprints, or reports from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
@@
After screening titles and abstracts, we shortlisted 250 papers. A subsequent full-text review narrowed the selection to 180 primary studies that directly addressed one or more TRiSM pillars. In addition to peer-reviewed literature, we also incorporated relevant technical reports, blogs, and website links to contextualize and map the emerging landscape of Agentic AI within the TRiSM framework. An analysis on the literature is shown in Figure 2.
3.4 Quality Assurance
To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [43, 42]. Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems? (2) Are the methods, experiments, or architectural designs well-documented and reproducible? (3) Does the paper substantively engage with at least one of the trust, risk, security, explainability, or privacy aspects? (4) Does the study offer empirical, theoretical, or normative contributions relevant to Agentic AI governance? Each criterion was rated as low, medium, or high. Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only. Quality ratings were assigned independently by four reviewers, with discrepancies resolved through discussion.
@@ 4 Fundamentals of Agentic AI Systems
In this section, we present the fundamentals of Agentic AI systems.
@@
Agentic AI systems fundamentally transform the design of autonomous systems, and are different from traditional AI agents. Traditional agents operate through predefined rules [44], heuristic workflows [45, 46], or deterministic logic [47], and perform well in narrow, well-defined environments. Cognitive Agents [48], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [49] or ACT-R [50]. However, they remain constrained by static knowledge bases and limited adaptability [48].
In contrast to traditional AI agents, Agentic AI systems leverage foundation models (mainly LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities. Agentic AI systems fundamentally redefine autonomy through three core innovations: (1) Multi-agent coordination: Specialized agents, such as planners and verifiers, collaborate using structured protocols [17, 19]. (2) Persistent context: Memory architectures maintain task state across workflows, ensuring continuity [51]. (3) Dynamic meta-orchestration: Systems delegate tasks and resolve conflicts dynamically [52].
This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [53]. For example, systems like AutoGen [18] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [54]. This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [19]. A comparison of traditional AI agents and Agentic AI systems is presented in Table 2.
@@
Autonomy Model × Reactive or deliberative (e.g., cognitive architectures) Goal-driven planning & adaptation
$\bullet$ Fixed action sequences or symbolic reasoning $\bullet$ Recursive self-improvement
Cognitive Foundation × Symbolic logic, FSMs, or cognitive models Foundation models (LLMs/LIMs)
$\bullet$ Hand-coded or structured KBs $\bullet$ Emergent reasoning
@@
Reasoning Approach × Deterministic or multi-step symbolic Multi-step & contextual
$\bullet$ Rule-based or production rules $\bullet$ Chain-of-Thought (CoT)
Collaboration × Isolated or limited coordination Role-specialized coordination
@@
Temporal Context × Episodic, often stateless Persistent memory
$\bullet$ Session-bound or episodic memory $\bullet$ VectorDB/LTM
Orchestration × Hard-coded or fixed workflows Dynamic meta-agents
$\bullet$ Sequential or predefined goals $\bullet$ Conflict resolution
Tool Utilization × Static or domain-specific tools Planned API invocation
@@
Context Awareness × Bounded or heuristic context Memory-augmented
$\bullet$ Limited or cognitive context $\bullet$ RAG architectures
Learning Mechanism × Limited or rule-based learning Emergent learning
$\bullet$ Fixed rules or chunking $\bullet$ Fine-tuning/RL
Human-AI Interaction × Scripted or cognitive modeling Natural language interfaces
$\bullet$ Limited or human-like reasoning $\bullet$ Conversational adaptability
Exemplars $\triangleright$ MYCIN  $\triangleright$ ELIZA  $\triangleright$ SOAR  $\triangleright$ ACT-R $\triangleright$ AutoGen  $\triangleright$ ChatDev  $\triangleright$ MetaGPT  $\triangleright$ AgentVerse
@@ 4.2 LLM-Based Agentic AI System Architecture
Agentic Multi-Agent System (AMAS) represent an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks. As illustrated in Figure 3, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem. At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation [55]. These agents access a Shared Toolchain Interface [56] to execute code, perform searches, or interact with domain-specific APIs.
@@
Figure 3: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms). The modular design highlights adaptive intelligence and inter-agent synergy.
Communication and coordination are facilitated through a Communication Middleware [57] that allows agents to share goals, observations, or intermediate results. A Task Manager or Orchestrator [58] governs high-level planning and delegates subtasks to agents based on their roles or specializations. Agents can read from and write to a World Model or Shared Memory [33], which stores contextual knowledge, system state, or evolving task data. Human oversight is supported through a Human-in-the-Loop Interface [46], enabling users to prompt, correct, or halt agent behavior.
To ensure accountability, a Trust and Audit module monitors agent actions, logs tool usage, and generates behavioral traces [59]. However, this modular and distributed structure introduces significant TRiSM [31] challenges. With multiple autonomous agents accessing external resources, the Security Gateway becomes critical for enforcing access controls, authentication, and sandboxing [60].
Likewise, a dedicated Privacy Management Layer is essential to prevent leakage of sensitive or personally identifiable information [61], especially when data traverses multiple agents or tools. Finally, an Explainability Interface must provide interpretable rationales for multi-agent decisions, supporting transparency and trust calibration [62]. Together, these architectural elements make AMAS powerful yet complex, raising unique and urgent questions about how to ensure their trustworthiness, mitigate systemic risks, and secure them against adversarial behaviors. Below, we discuss the architecture of AMAS.
@@
Language Model Core (Agent Brain). At the center of an Agentic AI system lies a LLM serving as the primary decision-making controller or “brain” [63]. The core LLM is initialized with a user goal and a structured agent prompt (defining its role, capabilities, and tool access). It then generates step-by-step decisions or actions, interpreting instructions, producing reasoning traces, and selecting next steps in either natural language or structured action formats. In many agent frameworks, such as AutoGPT [64], Baby AGI [65] and GPT Engineer [66], the LLM governs the full control loop, orchestrating the overall system behavior.
@@
Planning and Reasoning Module. To handle complex goals, an explicit planning mechanism decomposes tasks into manageable sub-goals. This can be done internally via chain-of-thought (CoT) or tree-of-thoughts prompting (ToT) [67], where the model performs intermediate reasoning before arriving at a final decision. Reasoning is also one aspect of cognitive functionalities [68]. Some implementations employ external planning systems by translating goals into structured planning languages and using classical planners for long-term decision-making. Planning is often interleaved with execution and feedback: the agent refines its plan based on outcomes, alternating between reasoning, acting, and integrating observations. Techniques like Reasoning and Acting (ReAct) [69] exemplify this loop by guiding the agent through repeated reasoning–action–observation cycles, improving performance on complex tasks.
@@
Memory Module. Agentic AI systems integrate memory to maintain context across iterations. This includes short-term memory (recent interactions held within the prompt context) and long-term memory (accumulated knowledge or experiences) [8]. Long-term memory is often implemented using vector databases, where key facts or past events are stored and retrieved by similarity search. By reintegrating past data into the LLM prompt, the agent can recall relevant information across sessions, avoid repetition, and support coherent long-term planning. Effective memory management enables adaptive, learning-driven behavior.
@@
Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes.
@@
AutoGPT [70] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
BabyAGI [65] GPT-3.5/4 Task queue In-memory Web search Minimal task generator
@@
GPT Engineer [71] GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
@@
LangGraph [72] Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
@@
AutoGen [18] GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
@@
MRKL [73] LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
@@
Reflexion [74] GPT-3.5/4 Retry–reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
@@
MetaGPT [75] GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
@@
Voyager [76] GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
@@
WebVoyager [77] LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
@@
HuggingGPT [78] ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
@@
CAMEL [79] GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
@@
ChatDev [80] GPT-4 Chat chain File repo Unix tools Simulated software development workflow
@@
CrewAI [81] Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
@@
AgentVerse [82] Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
@@
OpenAgents [83] Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
SuperAGI [84] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
Tool-Use Interface. To extend its capabilities beyond text generation, the agent is equipped with a tool-use interface [85]. This layer allows the invocation of external tools such as web search, APIs, code interpreters, or databases. The available tools are specified in the agent prompt with command schemas. When the LLM determines a tool is needed, it emits a structured command, which is executed externally. The result is fed back into the LLM as a new observation. This mechanism enables the agent to access real-time information, perform computations, and interact with external systems dynamically. Tool-use frameworks like Modular Reasoning, Knowledge, and Language (MRKL) [73] illustrate this design by routing queries to different expert modules (symbolic or learned) with an LLM as the router. Similarly, approaches like Toolformer [85] train the model to insert API calls in its text generation.
Perception and Environment Interface. For agents interacting with dynamic environments, such as web interfaces, simulated worlds, or physical systems, an observation-action interface is essential [86]. Perception modules translate raw inputs (e.g., sensor data, images, or textual states) into representations the LLM can process. Conversely, the agent’s chosen actions are executed within the environment, and the resulting state changes are returned to the agent as observations. This loop supports sense-plan-act cycles that continue until the task is completed or halted. In robotic or multimodal settings, the interface may include additional sensory models (e.g., vision transformers), but the core flow remains consistent.
Integration and Autonomy. These modules together form a closed-loop architecture. The LLM plans and reasons over tasks, guided by memory and tools, and interfaces with the environment to execute decisions and observe results [16]. Each iteration updates the agent’s context, enabling self-refinement through dynamic feedback. This integrated design empowers Agentic AI systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
Table  3 presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice. For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory to serve as both tool layer and environment interface.